Over the last three evenings, Heather received a total of 107 phone calls at the call center. The first evening, she received 8 fewer calls than the second evening. The third evening, she received 3 times as many calls as the second evening. How many phone calls did she receive each evening?
Answer #1

x be number of phone calls received on first evening y be number of phone calls received on second evening z be number of phox+y+z =100 -X&Y=8 2=34 2-3y=0

The coefficient determinant is D= The three determinants D1, D2 and D3 are obtained by replacing first, second and third colu